# Zewei Zhou Tianhui Cai11footnotemark: 1 Seth Z. Zhao  Yun Zhang  Zhiyu Huang Bolei Zhou  Jiaqi Ma

### v1

@@ 4.2 Main Results
RFT Performance. We apply RFT to the full-data CoT reasoning model trained via SFT to enhance its planning performance. As shown in Fig. 5(a), RFT yields a 10.6% improvement in PDMS (on the NAVSIM testing set) and a 66.8% reduction in runtime (average over 500 testing scenarios). The reward curve in Fig. 5(b) illustrates the progressive improvement of the model’s policy during RFT. Experiments with different GRPO group sample sizes indicate that larger groups lead to better performance by promoting broader exploration of training samples. As illustrated in Fig. 5(c), RFT also reduces unnecessary and slow reasoning in simple scenarios, driven by the CoT length penalty that encourages fast thinking for straightforward driving cases. A qualitative comparison shows that the SFT model produces suboptimal plans due to error accumulation in generation, whereas the RFT model (optimized via PDMS-based reward) generates better planning trajectories.
nuPlan Benchmark Results. We evaluate AutoVLA against state-of-the-art end-to-end driving models on the NAVSIM benchmark [51] and present results in Table 1. In best-of-N planning, we use an oracle scorer to select the optimal trajectory from six generated candidates. After RFT, AutoVLA demonstrates significantly improved performance, aligning more closely with the NAVSIM reward signal. The best-of-N strategy further enhances performance, achieving the highest PDMS. Overall, AutoVLA achieves competitive results while demonstrating scalability across diverse datasets.
@@
Waymo E2E Performance. We evaluate AutoVLA on the Waymo end-to-end driving dataset [52], which features long-tail and complex driving scenarios. The model’s performance under various training settings on the test set is presented in Fig. 6. The results reveal that pretraining on a combination of nuPlan and nuScenes datasets significantly enhances performance, suggesting enhanced scene understanding through exposure to more diverse training data. Incorporating CoT reasoning in training further improves planning performance compared to action-only training. Post-training with RFT, using ADE as the reward function, achieves the best overall RFS metric. A qualitative example in a construction zone demonstrates the model’s ability to reason about occlusions and generate effective detour plans. Additional results are provided in the Supplementary Material.
CARLA Closed-loop Performance. We evaluate the closed-loop driving performance of our AutoVLA model on the Bench2Drive benchmark [54] in the CARLA simulator. The model is trained using SFT with both trajectory-only and CoT data. During testing, the planning frequency is set to 2 Hz. The results, shown in Table 2, demonstrate that AutoVLA outperforms existing end-to-end driving models in terms of overall driving score and success rate in the closed-loop test.
@@
AD-MLP [99] 18.05 0.00 48.45 22.63
UniAD-Base [65] 45.81 16.36 129.21 43.58
VAD [66] 42.35 15.00 157.94 46.01
TCP-traj [100] 59.90 30.00 76.54 18.08
DriveAdapter [101] 64.22 33.08 70.22 16.01
Orion [42] 77.74 54.62 151.48 17.38
@@
4.3 Ablation Study
@@
PDM Score $\uparrow$ 59.24 80.54
Avg. L2 (m) $\downarrow$ 1.29 0.86
Avg. Col. (%) $\downarrow$ 0.98 0.35
Runtime (s) 9.31 3.95
Text Waypoint Output. We use the same mixed training set from the nuPlan and nuScenes datasets to train a model that predicts waypoints in a text format, which are then converted into a trajectory. We evaluate its performance in an open-loop planning setting using the standard test sets. The results, shown in Table 3, indicate that our action tokenization and generation method significantly outperforms the text-based waypoint prediction approach. Additionally, due to the need to decode numerical values, the text-based method incurs a substantially higher computational cost in generating the final trajectory. This shows the limitation of language models in handling precise numerical reasoning.
@@ 5 Conclusions
Limitation and Future Work. Although our model with dual-process adaptation achieves near-real-time inference (1 Hz), it remains highly GPU-dependent, requiring significant memory and computing. Future work will focus on real-time applications, optimizing runtime efficiency and reducing computation overhead (e.g., through model quantization) to enable real-time deployment.
@@ A.1 Action Codebook
To enable trajectory-level planning within a language model, we introduce physical action tokens, each representing a short-term feasible vehicle maneuver. These tokens are derived from clustering vehicle motion patterns in the Waymo Open Motion Dataset (WOMD) [102], which contains extensive real-world vehicle trajectories. The resulting codebook, visualized in Fig. S1(a), is used across experiments involving multiple real-world datasets, including nuPlan, nuScenes, and Waymo. Due to differences in vehicle dynamics compared to real-world data for simulation testing, we construct a separate action codebook using the same clustering procedure on the CARLA-Garage dataset [55]. This produces another set of 2048 action tokens for simulation testing, as shown in Fig. S1(b).
@@ A.2 Action Tokenizer
We implement an action tokenizer based on the constructed action codebook. During training, continuous trajectories are discretized by mapping each 0.5-second segment to its nearest action token in the codebook $\mathcal{A}$, resulting in a sequence of discrete tokens $[a_{1},a_{2},\dots,a_{T}]$. During inference, the language model autoregressively generates a sequence of action tokens, each representing a 0.5-second motion segment. These tokens are then converted back to their corresponding motion segments in the codebook and applied sequentially from the initial ego pose. Therefore, the model reconstructs a continuous trajectory in the ego-centric coordinate frame by composing these local displacements and rotations.
@@ Appendix B Reasoning Data Collection
A large-scale, high-quality reasoning dataset with chain-of-thought (CoT) annotations is essential for enabling robust reasoning capabilities in vision-language-action (VLA) models. In this paper, we introduce an automated reasoning annotation pipeline using the state-of-the-art Qwen2.5-VL-72B vision-language model [21], as illustrated in Fig. S2. The pipeline significantly reduces reliance on human annotations and facilitates effective knowledge distillation from a more powerful, large-scale model to a more efficient, compact model.
@@ B.1 Reasoning Annotation Pipeline
Reasoning Data Generation. We employ the most capable model in the Qwen-VL series as the reasoning data annotation engine, leveraging its strong reasoning capacity and extensive world knowledge. The maximum length for newly generated reasoning textual tokens is set to 700. Additionally, we extract answers to relevant questions from the DriveLM VQA dataset [96], which focuses on perception, prediction, and planning in the nuScenes and CARLA datasets, and then we reformat them into our standardized reasoning format. These reformatted samples are then combined with the generated data to construct the final reasoning dataset.
@@ B.2 Reasoning Annotation Visualization
Some CoT reasoning annotation examples from the Waymo end-to-end driving dataset are shown in Fig. S3, showcasing that our annotation pipeline can deal with situations where a vehicle already stopped at a stop sign can proceed, rather than remaining stopped indefinitely. The pipeline also accurately interprets construction-related road control scenarios, enabling high-quality reasoning annotations. Examples on the nuPlan dataset are provided in Fig. S4, demonstrating the pipeline’s capability to distinguish the functional relevance of stop signs and traffic lights across different lanes, rather than stopping at every detected stop sign.
@@ Appendix C Details of Supervised Fine-tuning
The model takes both system and user prompts as context inputs. As illustrated in Fig. S5, the system prompt defines the model’s role, task, and the expected CoT reasoning format. The user message describes the multi-view camera observations, the ego vehicle’s current state (i.e., speed and acceleration) and optionally historical states, and the high-level driving instruction. We use three camera views (front, front-left, and front-right), each providing four consecutive frames. Images in the video stream are resized to maintain original aspect ratios but reduced to $28\times 28\times 128$ pixels.
@@ Appendix D Details of Reinforcement Fine-tuning
Group Relative Policy Optimization (GRPO) employs group-based sampling to compute the advantage function, replacing conventional state-value estimators or critic models. This design accelerates training while aligning naturally with the inherent multimodality of planning, which requires evaluating and selecting from a set of candidate trajectories. The overall reinforcement fine-tuning (RFT) procedure is illustrated in Algorithm 1.
@@ D.1 Kullback-Leibler (KL) Divergence
The KL divergence term $\mathbb{D}_{\text{KL}}$ is incorporated in Eq. 3 to regularize the current policy with respect to the reference policy:
@@ D.2 Reward Function
The reward function provides the primary RFT training signal, guiding the optimization of policy updates. To further enhance comprehensive driving performance beyond imitating expert driving, we adopt a hybrid reward design that integrates the primary driving reward $r_{\text{Driving}}$ and a CoT penalty $r_{\text{CoT}}$ to discourage lengthy reasoning. For the NAVSIM Benchmark, we employ Predictive Driver Model Score (PDMS) [51] as the primary reward. For the Waymo end-to-end dataset, due to the limited annotation of RFS, we employ the (normalized) Average Displacement Error (ADE) as the driving reward. The final reward function, balanced by a weighting coefficient $\lambda_{r}$, is defined as:
@@
where the components include No at-fault Collision (NC), Drivable Area Compliance (DAC), Ego Progress (EP), Comfort (C), and Time-to-Collision (TTC). Detailed descriptions of each component can be found in [51]. Each sub-score is expressed as a percentage, and their weighted combination yields a normalized score within the range $[0,1]$, reflecting aspects of safety, comfort, and progress. For any scenario in which the planner fails (due to output errors), we assign a score of 00.
@@ D.3 Implementation Details
We use the navtrain split of the nuPlan dataset for RFT for the NAVSIM benchmark and the validation split of the Waymo E2E dataset for the Waymo benchmark. The vision encoder of pretrained AutoVLA is frozen, and the model is fine-tuned using Low-Rank Adaptation (LoRA) to reduce training costs and memory consumption. Specifically, both the LoRA rank and alpha are set to 8, with a dropout rate of 0.1. The pretrained SFT model is used as the reference policy during optimization. The hyperparameters $\gamma$, $L_{tol}$, and learning rate are set to $2\times 10^{-3}$, $400$, and $3\times 10^{-5}$, respectively. To ensure that the driving reward signal is dominant, the regularization weight is set to a relatively small value, $\lambda_{r}=0.3$. We also set $\delta=2$, $\kappa=10$ in the RFT of the Waymo dataset. Moreover, we configure the generation parameters with a sampling temperature of $1.0$, top-p of $1.0$, and top-k of $0.0$ to encourage diverse and exploratory generation during GRPO sampling, which effectively covers a wider range of possible actions.
@@ E.1 Data Preprocessing
To enable mixed training across multiple driving datasets, we develop a unified data preprocessing pipeline that standardizes the format across all datasets. For each sample, we extract and standardize: 1) Ground truth trajectory coordinates and headings in the ego vehicle’s coordinate frame at 2 Hz, 2) Image paths for multi-view camera image sequences consisting of 4 consecutive frames captured at 2 Hz (providing 2 seconds of history), 3) CoT reasoning annotations, 4) Vehicle states including its current velocity and acceleration, 5) High-level driving instructions. The preprocessing pipeline handles dataset-specific differences in data format, sampling rates, and coordinate systems to create a consistent format. The size and distribution of the final formatted dataset are shown in Table S1.
@@
Waymo. The Waymo E2E dataset provides 2037 training and 479 validation segments, each containing a 20-second video with driving logs for the entire duration. We sample reasoning data using a 4-second sliding window and extract trajectory-only data using a 2-second sliding window offset by 1 second from the reasoning samples. Due to noise in the position data, the estimated vehicle heading can exhibit abrupt fluctuations when the vehicle is stationary. To address this issue, we apply a motion threshold to detect stationary periods and smooth the heading accordingly. The test set comprises 1505 samples.
@@
nuPlan (NAVSIM) [51] 166.3k 45.6k 12.1k
nuScenes [53] 19.0k 2.9k 5.6k
Waymo [52] 23.8k 7.2k 1.5k
CARLA [96] 274.5k 53.2k

### v3

@@ 4.2 Main Results
RFT Performance. The slow-thinking mode incurs a significantly higher runtime due to the generation of CoT reasoning compared to the fast-thinking mode, as shown in Table˜2. To mitigate this overhead, we introduce RFT to enhance AutoVLA’s adaptive thinking capability and avoid unnecessary reasoning in straightforward scenarios. Specifically, we apply RFT to the full-data CoT reasoning model trained via SFT to enhance its planning performance. As shown in Fig.˜5(a), RFT yields a 10.6% improvement in PDMS (on the NAVSIM testing set) and a 66.8% reduction in runtime (average over 500 testing scenarios). The reward curve in Fig.˜5(b) illustrates the progressive improvement of the model’s policy during RFT. Experiments with different GRPO group sample sizes indicate that larger groups lead to better performance by promoting broader exploration of training samples. As illustrated in Fig.˜5(c), RFT also reduces unnecessary and slow reasoning in simple scenarios, driven by the CoT length penalty that encourages fast thinking for straightforward driving cases. A qualitative comparison shows that the SFT model produces suboptimal plans due to error accumulation in generation, whereas the RFT model (optimized via PDMS-based reward) generates better trajectories.
nuPlan Benchmark Results. We evaluate AutoVLA against state-of-the-art end-to-end driving models on the NAVSIM benchmark [dauner2024navsim] and present results in Table˜1. In best-of-N planning, we use an oracle scorer to select the optimal trajectory from six generated candidates. After RFT, AutoVLA demonstrates significantly improved performance, aligning more closely with the NAVSIM reward signal. The best-of-N strategy further enhances performance, achieving the highest PDMS. Overall, AutoVLA achieves competitive results while demonstrating scalability across diverse datasets.
@@
Waymo E2E Performance. We evaluate AutoVLA on the Waymo end-to-end driving dataset [xu2025wod], which features long-tail and complex driving scenarios. The model’s performance under various training settings on the test set is presented in Fig.˜6. The results reveal that pretraining on a combination of nuPlan and nuScenes datasets significantly enhances performance, suggesting enhanced scene understanding through exposure to more diverse training data. Incorporating CoT reasoning in training further improves planning performance compared to action-only training. Post-training with RFT, using ADE as the reward function, achieves the best overall RFS metric. A qualitative example in a construction zone demonstrates the model’s ability to reason about occlusions and generate effective detour plans.
CARLA Closed-loop Performance. We evaluate the closed-loop driving performance of our AutoVLA model on the Bench2Drive benchmark [jia2024bench2drive] in the CARLA simulator. The model is trained using SFT with both trajectory-only and CoT data. During testing, the planning frequency is set to 2 Hz. The results, shown in Table˜3, demonstrate that AutoVLA outperforms existing end-to-end driving models in terms of overall driving score and success rate in the closed-loop test.
@@
AD-MLP [zhai2023rethinking] 18.05 0.00 48.45 22.63
UniAD-Base [hu2023planning] 45.81 16.36 129.21 43.58
VAD [jiang2023vad] 42.35 15.00 157.94 46.01
TCP-traj [wu2022trajectory] 59.90 30.00 76.54 18.08
DriveAdapter [jia2023driveadapter] 64.22 33.08 70.22 16.01
Orion [fu2025orion] 77.74 54.62 151.48 17.38
@@
4.3 Ablation Studies
Table 4: Action Tokenization Accuracy with Different Codebook Sizes and Methods. (DCT: discrete cosine transform; MC: movement coverage; CU: codebook usage. Bold indicates best performance, underline the second-best, and tilde the third-best.)
Codebook Size RT-1 (Action Bin) [brohan2022rt] FAST (DCT) [pertsch2025fast] K-disk (Ours)
ADE(m) $\downarrow$ FDE(m) $\downarrow$ ADE(m) $\downarrow$ FDE(m) $\downarrow$ ADE(m) $\downarrow$ FDE(m) $\downarrow$ MC $\uparrow$ CU $\uparrow$
256 0.1440 0.2942 0.1708 0.2137 0.0687 0.1034 86.47% 100.0%
1024 0.1052 0.1883 0.0522 0.0588 0.0253 0.0282 97.41% 100.0%
2048 0.1014 0.1775 0.0281 0.0309 0.0182 0.0203 99.42% 100.0%
4096 0.1001 0.1739 0.0149 0.0161 0.0141 0.0155 100.0% 91.46%
Text Waypoint Output. We use the same mixed training set from the nuPlan and nuScenes datasets to train a model that predicts waypoints in a text format, which are then converted into a trajectory of waypoints. We evaluate its performance in an open-loop planning setting using the standard test sets. The results, shown in Table˜5, indicate that our action tokenization and generation method significantly outperforms the text-based waypoint prediction approach. Additionally, due to the need to decode
@@ 4.3 Ablation Studies
PDM Score $\uparrow$ 71.31 80.54
Avg. L2 (m) $\downarrow$ 0.89 0.70
Avg. Col. (%) $\downarrow$ 0.36 0.31
Runtime (s) $\downarrow$ 7.65 3.95
Table 6: Performance Comparison across Different Tokenization Methods on NAVSIM
Tokenization PDMS $\uparrow$ Collision $\uparrow$ Progress $\uparrow$
FAST (DCT) [pertsch2025fast] 67.63 92.74 64.09
K-disk (Ours) 80.54 96.89 75.82
numerical values, the text-based method incurs a substantially higher computational cost in generating the final trajectory. This shows the limitation of language models in handling precise numerical reasoning.
Action Tokenization Methods. We evaluate reconstruction accuracy across different codebook sizes for several tokenization methods, including RT-1 [brohan2022rt], FAST [pertsch2025fast], and our proposed K-disk tokenization, as shown in Table˜4. For RT-1, acceleration and steering rate are discretized using uniform action bins, and a kinematic model is employed to reconstruct the trajectory. However, because only trajectory-level data is available and control actions must be indirectly inferred, this binning approach yields the highest reconstruction error. The FAST tokenization method is more sensitive to codebook size than our K-disk approach, achieving comparable reconstruction accuracy only at $K=4096$. In contrast, our proposed method consistently attains the highest reconstruction accuracy across all codebook sizes.
When the codebook size is small ($K=256$), reconstruction errors are significantly higher due to limited movement coverage, as many movements cannot be adequately represented by the small set of codebook tokens. As the codebook size increases beyond $K=2048$, the improvements in tokenization accuracy and movement coverage become marginal, while codebook usage decreases as many tokens remain redundant and unused. Considering this trade-off, we select $K=2048$ to balance reconstruction accuracy, movement coverage, and efficient codebook usage.
Furthermore, Table˜6 reports planning performance on NAVSIM (nuPlan) for different tokenization methods. Our proposed method outperforms FAST. In FAST, the discrete cosine transform converts fixed-horizon planning trajectories into variable-length token sequences, complicating the selection of appropriate token lengths and making fixed-horizon planning difficult.
@@ 5 Conclusions
Limitation and Future Work. Although our model with dual-process adaptation achieves near-real-time inference (1 Hz), it remains highly GPU-dependent, requiring significant memory and computing. Future work will focus on real-time applications, optimizing runtime efficiency, and reducing computation overhead (e.g., through model quantization) to enable real-time deployment.
Acknowledgements
This work was supported by the Federal Highway Administration Center of Excellence on New Mobility and Automated Vehicles, and by the National Science Foundation under Award No. 2346267, POSE: Phase II - DriveX: An Open Source Ecosystem for Automated Driving and Intelligent Transportation Research.
[Uncaptioned image]AutoVLA Supplementary Material
@@ A.1 Action Codebook
To enable trajectory-level planning within a language model, we introduce physical action tokens, each representing a short-term feasible vehicle maneuver. These tokens are derived from clustering vehicle motion patterns in the Waymo Open Motion Dataset (WOMD) [ettinger2021large], which contains extensive real-world vehicle trajectories. The resulting codebook, visualized in Fig.˜S1(a), is used across experiments involving multiple real-world datasets, including nuPlan, nuScenes, and Waymo. Due to differences in vehicle dynamics compared to real-world data for simulation testing, we construct a separate action codebook using the same clustering procedure on the CARLA-Garage dataset [Jaeger_2023_ICCV]. This produces another set of 2048 action tokens for simulation testing, as shown in Fig.˜S1(b).
@@ A.2 Action Tokenizer
We implement an action tokenizer based on the constructed action codebook. During training, continuous trajectories are discretized by mapping each 0.5-second segment to its nearest action token in the codebook $\mathcal{A}$, resulting in a sequence of discrete tokens $a_{1},a_{2},\dots,a_{T}]$. During inference, the language model autoregressively generates a sequence of action tokens, each representing a 0.5-second motion segment. These tokens are then converted back to their corresponding motion segments in the codebook and applied sequentially from the initial ego pose. Therefore, the model reconstructs a continuous trajectory in the ego-centric coordinate frame by composing these local displacements and rotations.
@@ Appendix B Reasoning Data Collection
A large-scale, high-quality reasoning dataset with chain-of-thought (CoT) annotations is essential for enabling robust reasoning capabilities in vision-language-action (VLA) models. In this paper, we introduce an automated reasoning annotation pipeline using the state-of-the-art Qwen2.5-VL-72B vision-language model [bai2025qwen2], as illustrated in Fig.˜S2. The pipeline significantly reduces reliance on human annotations and facilitates effective knowledge distillation from a more powerful, large-scale model to a more efficient, compact model.
@@ B.1 Reasoning Annotation Pipeline
Reasoning Data Generation. We employ the most capable model in the Qwen-VL series as the reasoning data annotation engine, leveraging its strong reasoning capacity and extensive world knowledge. The maximum length for newly generated reasoning textual tokens is set to 700. Additionally, we extract answers to relevant questions from the DriveLM VQA dataset [sima2024drivelm], which focuses on perception, prediction, and planning in the nuScenes and CARLA datasets, and then we reformat them into our standardized reasoning format. These reformatted samples are then combined with the generated data to construct the final reasoning dataset.
@@ B.2 Reasoning Annotation Visualization
Some CoT reasoning annotation examples from the Waymo end-to-end driving dataset are shown in Fig.˜S3, showcasing that our annotation pipeline can deal with situations where a vehicle already stopped at a stop sign can proceed, rather than remaining stopped indefinitely. The pipeline also accurately interprets construction-related road control scenarios, enabling high-quality reasoning annotations. Examples on the nuPlan dataset are provided in Fig.˜S4, demonstrating the pipeline’s capability to distinguish the functional relevance of stop signs and traffic lights across different lanes, rather than stopping at every detected stop sign.
@@ Appendix C Details of Supervised Fine-tuning
The model takes both system and user prompts as context inputs. As illustrated in Fig.˜S5, the system prompt defines the model’s role, task, and the expected CoT reasoning format. The user message describes the multi-view camera observations, the ego vehicle’s current state (i.e., speed and acceleration) and optionally historical states, and the high-level driving instruction. We use three camera views (front, front-left, and front-right), each providing four consecutive frames. Images in the video stream are resized to maintain original aspect ratios but reduced to $28\times 28\times 128$ pixels.
@@ Appendix D Details of Reinforcement Fine-tuning
Group Relative Policy Optimization (GRPO) employs group-based sampling to compute the advantage function, replacing conventional state-value estimators or critic models. This design accelerates training while aligning naturally with the inherent multimodality of planning, which requires evaluating and selecting from a set of candidate trajectories. The overall reinforcement fine-tuning (RFT) procedure is illustrated in Algorithm˜1.
@@ D.1 Kullback-Leibler (KL) Divergence
The KL divergence term $\mathbb{D}_{\text{KL}}$ is incorporated in Eq.˜3 to regularize the current policy with respect to the reference policy:
@@ D.2 Reward Function
The reward function provides the primary RFT training signal, guiding the optimization of policy updates. To further enhance comprehensive driving performance beyond imitating expert driving, we adopt a hybrid reward design that integrates the primary driving reward $r_{\text{Driving}}$ and a CoT penalty $r_{\text{CoT}}$ to discourage lengthy reasoning. For the NAVSIM Benchmark, we employ Predictive Driver Model Score (PDMS) [dauner2024navsim] as the primary reward. For the Waymo end-to-end dataset, due to the limited annotation of RFS, we employ the (normalized) Average Displacement Error (ADE) as the driving reward. The final reward function, balanced by a weighting coefficient $\lambda_{r}$, is defined as:
@@
where the components include No at-fault Collision (NC), Drivable Area Compliance (DAC), Ego Progress (EP), Comfort (C), and Time-to-Collision (TTC). Detailed descriptions of each component can be found in [dauner2024navsim]. Each sub-score is expressed as a percentage, and their weighted combination yields a normalized score within the range $[0,1]$, reflecting aspects of safety, comfort, and progress. For any scenario in which the planner fails (due to output errors), we assign a score of 0.
@@ D.3 Implementation Details
We use the navtrain split of the nuPlan dataset for RFT for the NAVSIM benchmark and the validation split of the Waymo end-to-end driving dataset for the Waymo benchmark. The vision encoder of pretrained AutoVLA is frozen, and the model is fine-tuned using Low-Rank Adaptation (LoRA) to reduce training costs and memory consumption. Specifically, both the LoRA rank and alpha are set to 8, with a dropout rate of 0.1. The pretrained SFT model is used as the reference policy during optimization. The hyperparameters $\gamma$, $L_{tol}$, and learning rate are set to $2\times 10^{-3}$, $400$, and $3\times 10^{-5}$, respectively. To ensure that the driving reward signal is dominant, the regularization weight is set to a relatively small value, $\lambda_{r}=0.3$. We also set $\delta=2$, $\kappa=10$ in the RFT of the Waymo dataset. Moreover, we configure the generation parameters with a sampling temperature of $1.0$, top-p of $1.0$, and top-k of $0.0$ to encourage diverse and exploratory generation during GRPO sampling, which effectively covers a wider range of possible actions.
@@ E.1 Data Preprocessing
To enable mixed training across multiple driving datasets, we develop a unified data preprocessing pipeline that standardizes the format across all datasets. For each sample, we extract and standardize: 1) Ground truth trajectory coordinates and headings in the ego vehicle’s coordinate frame at 2 Hz, 2) Image paths for multi-view camera image sequences consisting of 4 consecutive frames captured at 2 Hz (providing 2 seconds of history), 3) CoT reasoning annotations, 4) Vehicle states including its current velocity and acceleration, 5) High-level driving instructions. The preprocessing pipeline handles dataset-specific differences in data format, sampling rates, and coordinate systems to create a consistent format. The size and distribution of the final formatted dataset are shown in Table˜S1.
@@
Waymo. The Waymo end-to-end driving dataset provides 2037 training and 479 validation segments, each containing a 20-second video with driving logs for the entire duration. We sample reasoning data using a 4-second sliding window and extract trajectory-only data using a 2-second sliding window offset by 1 second from the reasoning samples. Due to noise in the position data, the estimated vehicle heading can exhibit abrupt fluctuations when the vehicle is stationary. To address this issue, we apply a motion threshold to detect stationary periods and smooth the heading accordingly. The test set comprises 1505 samples.
@@
nuPlan (NAVSIM) [dauner2024navsim] 166.3k 45.6k 12.1k
nuScenes [caesar2020nuscenes] 19.0k 2.9k 5.6k
Waymo [xu2025wod] 23.8k 7.2k 1.5k
CARLA [sima2024drivelm] 274.5k 53.2k